please help...
1. a triangle has a perimeter of 69 inches. the second side is 7 inches longer than the first side. the third side is 6 inches shorter than twice the length of the first side. find the length of each side.
2. michelle collected $184 selling girl scout cookies. the money consisted of $1, $5, and $10 bills. there were three times as many $1 bills as $5. if there were 60 bills, how many of each bill did she have?

Answer by scott8148(6628)

1. let x equal the first side, so x+7 equals the second side and 2x-6 equals the third side
x+(x+7)+(2x-6)=69 ... 4x+1=69 ... x=17 ...x+7=24 ...2x-6=28

2. let x equal $5 bills, so 3x equals $1 bills and 60-4x equals $10 bills
(5*x)+(1*3x)+(10*(60-4x))=184 ... 5x+3x+600-40x=184 ... -32x=-416 ... x=13 ... 3x=39 ... 60-4x=8